# Position: The Current AI Conference Model is Unsustainable! Diagnosing the Crisis of Centralized AI Conference

### v3

@@ Abstract
11footnotetext: nuochen@comp.nus.edu.sg
@@ 1 Introduction
Academic conferences are the lifeblood of the Artificial Intelligence (AI) community. They serve as primary venues for disseminating cutting-edge research, facilitating invaluable networking, and shaping the trajectory of the field. The growth of major conferences like NeurIPS, ICML, and ICLR is a testament to the dynamism and impact of AI research. Their core mission can be distilled into four pillars: (1) The Scientific Mission: Promoting AI research and enhancing academic exchange by serving as an efficient, peer-reviewed knowledge exchange platform. (2) Knowledge Dissemination: Sharing research results and recognizing thought leaders through presentations and awards. (3) Community Building: Fostering collaborations and a sense of belonging among researchers. (4) The Social Contract: Promoting diversity, equity, and inclusion (DEI) through inclusive practices. However, this success has come at a cost. The prevailing model—a handful of massive, centralized, in-person events—is straining under its own weight 1, 2, raising a critical question: Is the current AI conference model sustainable?
The challenges are multifaceted: (1) The rapid growth in paper submissions (see Figure 1) has overwhelmed the high-pressured peer-review system, often compromising review quality and fostering a culture where constructive criticism is discouraged and creates a hyper-competitive atmosphere that can stifle creativity; (2) Environmentally, the carbon footprint (see Figure 5) generated by thousands of researchers traveling across the globe is significant and increasingly at odds with global sustainability goals 3, 4; (3) The high-stakes environment exacts a human toll. Researchers, especially doctoral students, face immense pressure (see Figure 6) to publish in these top-tier venues, leading to widespread anxiety, job insecurity, and a diminished sense of academic belonging 5; (4) The sheer scale of these events not only makes meaningful connections difficult, ironically isolating individuals within a crowd of thousands, but also creates venue capacity bottlenecks (see Figure 8). The sheer scale of these events can also make meaningful connection difficult, ironically isolating individuals within a crowd of thousands.
@@
In this position paper, we argue that the current centralized AI conference model is increasingly unsustainable, as it undermines the core goals of such conferences (detailed in Table 1). Drawing from prior analyses 6, 7, we identify four primary goals:
@@
To address these, we propose the Community-Federated Conference (CFC) model—a decentralized framework that decouples peer review, dissemination, and networking into interconnected layers: unified global peer-review with rolling cycles, federated regional hubs for localized interactions, and a digital layer for global connectivity. This model mitigates the identified challenges, aligning better with the four core goals. We conclude with a call for the AI community to embrace this shift toward sustainability and equity.
@@ 2 Goals and Core Elements of AI Conferences
Artificial intelligence academic conferences (e.g., NeurIPS 8, ICLR 9, and ICML 10) have evolved from mere platforms for the release of technical achievements into multifunctional ecosystems with far-reaching influence, undertaking multiple missions such as promoting industry development, interdisciplinary integration, and social responsibility. In this section, we analyse the goals and core elements of AI conferences as following:
@@
The most basic and core goal of AI conferences is to serve as a efficient, peer-reviewed knowledge exchange platform 7 to prompt AI research 6, 11. For instance, ICML explicitly calls for papers presenting “significant, original, and previously unpublished research” in all areas of machine learning 10, highlighting its commitment to advancing the scientific frontier. Similarly, ICLR highlights its interest in “cutting-edge research on all aspects of deep learning, representation learning, and related areas”9, positioning itself as a primary venue for presenting state-of-the-art discoveries. Being accepted by a top AI conference means that a research has been recognized by the community and valued its exchange for advancing the AI domain.
@@ 4 Challenges in Current AI Conferences
While the core goals of AI conferences remain essential, the current model increasingly deviates from these objectives due to unsustainable growth in submission volumes. This explosive expansion not only strains resources but also undermines the missions outlined in Section 2, transforming conferences from focused knowledge hubs into overwhelming, high-volume events that prioritize quantity over depth (Mamba 21’s reject as a case). To illustrate, we analyze publication trends across computer science subfields using conferences categorization (see Table in the appendix) and conference data in CSRankings 16. As shown in Figure 1, while fields like Theory, Systems, and Interdisciplinary areas exhibit relatively stable growth over the past decade, AI publications have experienced an almost exponential surge ($R^{2}=0.979,p<0.001$ under a log-linear regression model) over the past decade (especially true for 2023 and 2024, likely driven by the emergence and proliferation of LLMs). This trend is estimated to increase to over 65,000 (95% prediction lower bound is 66,572) by 2030, more than tripling the 2024 figure of 21,400. Such unchecked growth has sparked widespread concerns such as the integrity of peer-review systems 1, 2. This hyper-competitive environment, where the sheer volume of submissions overwhelms the review process, may also lower the barrier for academic misconduct, such as abuse of LLM writing 22 and LLM-based prompt injection 23 in paper submissions, or make it harder to detect. Moreover, the annual cycle of conference submissions struggles to keep pace with the shrinking lifecycle of AI research. It is now a common phenomenon for work to be outdated by the time it is presented, and an initial rejection often leaves a paper with insufficient novelty for subsequent acceptance.
@@
This surge in paper volumes also exacerbates a cascade of interconnected issues for authors, reviewers, and organizers, including elevated carbon footprints from increased travel and computing demands. Since most AI conferences 8, 9, 10 mandate in-person attendance for accepted authors, publication growth generates massive travel volumes. For the largest AI conference NeurIPS 2024, we estimate emissions for 3,836 unique first authors (from 4,037 before de-duplication, with 13,307 actual attendees 17) flying round-trip from their organizations—concentrated in Asia, Europe, and the Americas (as highlighted in Figure 4)—to the Vancouver Convention Centre in Canada. Using the formula Activity data $\times$ Emission factors = emissions (details in the appendix), flight emissions alone reach 8,254 tCO2e, exceeding the daily carbon output of Vancouver’s approximately 680 thousand residents (based on 2.5 MtCO2e annually 20, or roughly 6,849 tCO2e per day). These environmental impacts can challenge conferences’ DEI initiatives, making it harder for researchers from underrepresented regions or those who are eco-conscious to participate. They also place unnecessary financial and time burdens on authors, including costs for visas and airfare.
@@
Prior studies have roughly examined travel impacts in non-computing conferences 24, 4, and provided services for carbon footprint calculations 3, 4. To extend the analysis to AI conference, we evaluate recent emissions from ICML (a 5-day event, based on top 200 first-author organizations), focusing on travel (transportation and accommodation) and venue (food and energy consumption); ICLR details appear in the appendix. As shown in Figure 5, transportation dominates emissions for non-hybrid ICML and ICLR, aligning with proportions from 25, while emissions from accepted authors have exploded—more than quadrupling over the past five years.
@@
The growing scale and competitive nature of AI conferences also affect the emotional well-being and sentiment of different participant groups (authors, reviewers, attendees). For example, 5 analyzed the academic impact on high pressure and job insecurity, and 26 raised that doctoral students commonly experience identity anxiety and a lack of academic belonging. To systematically investigate these sentiments within the community, we analyzed discussions on public forums. We queried the Reddit subreddit, r/MachineLearning, using terms such as ICLR, NeurIPS, ICML, and “machine learning conferences” to identify the top 25 of the most relevant and high-activity (labelled “hot”) discussion threads. We analyzed comments of each thread using VADER (Valence Aware Dictionary and sEntiment Reasoner), a lexicon-based sentiment analysis tool optimized for social media text. We focused on detecting indicators of negative sentiment or suggestions for improvement, and the sampled representative comments were subsequently aggregated to construct a sentence cloud in Figure 8, highlighting recurring themes and concerns within the community and the high overlap with bad issues in this paper (see supplementary material for the sources).
@@
Authors’ psychological stress may also stem from related dynamics, as evidenced by NeurIPS submission statistics in Figure 8. Under the pressure of massive submissions, accepted papers exhibit roughly linear growth ($R^{2}=0.964,p<0.001$ under a linear regression model,), while rejections surge at a much faster, near-exponential rate ($R^{2}=0.937$, $p<0.001$ under a log-linear regression model). This inevitably drives more resubmissions, further straining reviewer capacity. Compounding the issue, state-of-the-art (SOTA) benchmarks from prior submissions may become outdated by the next cycle, fostering a distorted mindset among authors who chase overly positive reviews at the expense of genuine innovation. For instance, studies suggest that AI agent capabilities double approximately every seven months 27, 28. Since the conference cycle from submission to presentation also lasts nearly seven months, this means research can be outdated by the time it is published, rendering a significant portion of the community’s effort inefficient. Such dynamics undermine community building and knowledge sharing. Additionally, the sheer volume of acceptances burdens organizers, with exploding attendance, visa restrictions, and other logistical constraints prompting NeurIPS to adopt a hybrid online-offline format since 2022. Pushing this concept further, NeurIPS 2025 will host parallel main conferences in Mexico City 29 and Copenhagen 30. These venue-related challenges are discussed below.
@@
As academic conferences grow in scale and scope, physical venues are increasingly unable to keep pace. This pressure is particularly evident at flagship AI conferences such as NeurIPS. The Vancouver Convention Centre, host of NeurIPS 2024, has a maximum capacity of approximately 18,000 attendees 31. Recognizing that registration demand was nearing this limit (as reflected in Figure 8), NeurIPS 2024 implemented a lottery system 32 for non-author registrations. While necessary from a logistical perspective, this policy introduced a layer of artificial scarcity that limits participation for students, early-career researchers, and unaffiliated attendees who may benefit most from in-person engagement.
@@ 5 Evaluating Incremental Fixes: Why Palliative Measures Fail
A case in point is the proposal to limit submissions per author 33 applied to manage submission volume. However, this supply-side constraint is the zero-sum game of submission caps. It does not reduce the immense institutional pressure to publish; it merely shifts it, forcing researchers to be more strategic but no less stressed. Such caps can disproportionately penalize junior researchers who need to build their publication record, or productive labs working on multiple innovative fronts. Most importantly, it does nothing to alleviate the systemic “publish-or-perish” culture that drives burnout and discourages high-risk, long-term research. It’s an administrative patch on a cultural wound.
Similarly, the adoption of multi-site or satellite conferences, such as NeurIPS 2025, hosting parallel events in Mexico City and Copenhagen 29, 30, is a direct response to venue capacity limits and aims to reduce long-haul travel for some participants. However, a multi-site conference still concentrates the entire community’s review burden and mental health anxiety into the same frantic, annual cycles. It maintains the centralized authority and the high-stakes, all-or-nothing evaluation process. It mitigates some travel but still fails to address the core issues of reviewer burnout, escalating author effort, and the exclusion inherent in any time-bound, high-cost event. It may also inadvertently create a two-tiered system, where one venue is perceived as more prestigious, thus contradicting the goal of equitable participation under the social contract.
@@ Appendix A Detailed Emission Factors
Air Travel 0.158 t CO2e / passenger.km 34
Accommodation (Kigali, 2023) 28.0 kg CO2e / room-night 35, 19
Accommodation (Vienna, 2024) 13.9 kg CO2e / room-night 36, 37
Accommodation (Singapore, 2025) 24.5 kg CO2e / room-night 38
Catering (Meals) 1.88 kg CO2e / meal 39, 40
Venue Operations 2.896 kg CO2e / person/day 41
The data in table 2 is retrieved from the UK’s Department for Environment, Food and Rural Affairs (DEFRA) GreenHouse Gas (GHG) Conversion Factors 202542.

### v4

@@ Abstract
footnotetext: nuochen@comp.nus.edu.sg
@@ 1 Introduction
Academic conferences are the lifeblood of the Artificial Intelligence (AI) community. They serve as primary venues for disseminating cutting-edge research, facilitating invaluable networking, and shaping the trajectory of the field. The growth of major conferences like NeurIPS, ICML, and ICLR is a testament to the dynamism and impact of AI research. Their core mission can be distilled into four pillars: (1) The Scientific Mission: Promoting AI research and enhancing academic exchange by serving as an efficient, peer-reviewed knowledge exchange platform. (2) Knowledge Dissemination: Sharing research results and recognizing thought leaders through presentations and awards. (3) Community Building: Fostering collaborations and a sense of belonging among researchers. (4) The Social Contract: Promoting diversity, equity, and inclusion (DEI) through inclusive practices. However, this success has come at a cost. The prevailing model (a handful of massive, centralized, in-person events) is straining under its own weight 1, 2, raising a critical question: Is the current AI conference model sustainable?
The challenges are multifaceted: (1) The rapid growth in paper submissions (see Figure 1) has overwhelmed the high-pressured peer-review system, often compromising review quality and fostering a culture where constructive criticism is discouraged and creates a hyper-competitive atmosphere that can stifle creativity; (2) Environmentally, the carbon footprint (see Figure 5) generated by thousands of researchers traveling across the globe is significant and increasingly at odds with global sustainability goals 3, 4; (3) The high-stakes environment exacts a human toll. Researchers, especially doctoral students, face immense pressure (see Figure 6) to publish in these top-tier venues, leading to widespread anxiety, job insecurity, and a diminished sense of academic belonging 5; (4) The sheer scale of these events not only makes meaningful connections difficult, ironically isolating individuals within a crowd of thousands, but also creates venue capacity bottlenecks (see Figure 8).
@@
In this position paper, we argue that the current centralized AI conference model is increasingly unsustainable, as it undermines the core goals of such conferences. Drawing from prior analyses 6, 7, we identify four primary goals detailed in Table 1.
@@
To address these, we propose the Community-Federated Conference (CFC) model, a decentralized framework that decouples peer review, dissemination, and networking into interconnected layers: unified global peer-review with rolling cycles, federated regional hubs for localized interactions, and a digital layer for global connectivity. This model mitigates the identified challenges, aligning better with the four core goals. We conclude with a call for the AI community to embrace this shift toward sustainability and equity.
@@ 2 Goals and Core Elements of AI Conferences
Artificial intelligence (AI) academic conferences (e.g., NeurIPS 8, ICLR 9, and ICML 10) have evolved from mere platforms for the release of technical achievements into multifunctional ecosystems with far-reaching influence, undertaking multiple missions such as promoting industry development, interdisciplinary integration, and social responsibility. In this section, we analyze the goals and core elements of AI conferences as follows:
@@
The most basic and core goal of AI conferences is to serve as an efficient, peer-reviewed knowledge exchange platform 7 to promote AI research 6, 11. For instance, ICML explicitly calls for papers presenting “significant, original, and previously unpublished research” in all areas of machine learning 10, highlighting its commitment to advancing the scientific frontier. Similarly, ICLR highlights its interest in “cutting-edge research on all aspects of deep learning, representation learning, and related areas”9, positioning itself as a primary venue for presenting state-of-the-art discoveries. Being accepted by a top AI conference means that a research has been recognized by the community and valued its exchange for advancing the AI domain.
@@ 4 Challenges in Current AI Conferences
While the core goals of AI conferences remain essential, the current model increasingly deviates from these objectives due to unsustainable growth in submission volumes. This explosive expansion not only strains resources but also undermines the missions outlined in Section 2, transforming conferences from focused knowledge hubs into overwhelming, high-volume events that prioritize quantity over depth (Mamba 21’s rejection as a case). To illustrate, we analyze publication trends across computer science subfields using conferences categorization (see Table in the appendix) and conference data in CSRankings 16. As shown in Figure 1, while fields like Theory, Systems, and Interdisciplinary areas exhibit relatively stable growth over the past decade, AI publications have experienced an almost exponential surge ($R^{2}=0.979,p<0.001$ under a log-linear regression model) over the past decade (especially true for 2023 and 2024, likely driven by the emergence and proliferation of LLMs). This trend is estimated to increase to over 65,000 (95% prediction lower bound is 66,572) by 2030, more than tripling the 2024 figure of 21,400. Such unchecked growth has sparked widespread concerns such as the integrity of peer-review systems 22, 1, 2. This hyper-competitive environment, where the sheer volume of submissions overwhelms the review process, may also lower the barrier for academic misconduct, such as abuse of LLM writing 23 and LLM-based prompt injection 24 in paper submissions, or make it harder to detect. Moreover, the annual cycle of conference submissions struggles to keep pace with the shrinking lifecycle of AI research. It is now a common phenomenon for work to be outdated by the time it is presented, and an initial rejection often leaves a paper with insufficient novelty for subsequent acceptance.
@@
This surge in paper volumes also exacerbates a cascade of interconnected issues for authors, reviewers, and organizers, including elevated carbon footprints from increased travel and computing demands. Since most AI conferences 8, 9, 10 mandate in-person attendance for accepted authors, publication growth generates massive travel volumes. For the largest AI conference NeurIPS 2024, we estimate emissions for 3,836 unique first authors (from 4,037 before de-duplication, with 13,307 actual attendees 17) flying round-trip from their organizations, concentrated in Asia, Europe, and the Americas (as highlighted in Figure 4), to the Vancouver Convention Centre in Canada. Using the formula Activity data $\times$ Emission factors = emissions (details in the appendix), flight emissions alone reach 8,254 tCO2e, exceeding the daily carbon output of Vancouver’s approximately 680 thousand residents (based on 2.5 MtCO2e annually 20, or roughly 6,849 tCO2e per day). These environmental impacts can challenge conferences’ DEI initiatives, making it harder for researchers from underrepresented regions or those who are eco-conscious to participate. They also place unnecessary financial and time burdens on authors, including costs for visas and airfare.
@@
Prior studies have roughly examined travel impacts in non-computing conferences 25, 4, and provided services for carbon footprint calculations 3, 4. To extend the analysis to AI conference, we evaluate recent emissions from ICML (a 5-day event, based on top 200 first-author organizations), focusing on travel (transportation and accommodation) and venue (food and energy consumption); ICLR details appear in the appendix. As shown in Figure 5, transportation dominates emissions for non-hybrid ICML and ICLR, aligning with proportions from 26, while emissions from accepted authors have exploded, more than quadrupling over the past five years.
@@
The growing scale and competitive nature of AI conferences also affect the emotional well-being and sentiment of different participant groups (authors, reviewers, attendees). For example, 5 analyzed the academic impact on high pressure and job insecurity, and 27 raised that doctoral students commonly experience identity anxiety and a lack of academic belonging. To systematically investigate these sentiments within the community, we analyzed discussions on public forums. We queried the Reddit subreddit, r/MachineLearning, using terms such as ICLR, NeurIPS, ICML, and “machine learning conferences” to identify the top 25 of the most relevant and high-activity (labelled “hot”) discussion threads. We analyzed comments of each thread using VADER (Valence Aware Dictionary and sEntiment Reasoner), a lexicon-based sentiment analysis tool optimized for social media text. We focused on detecting indicators of negative sentiment or suggestions for improvement, and the sampled representative comments were subsequently aggregated to construct a sentence cloud in Figure 8, highlighting recurring themes and concerns within the community and the high overlap with bad issues in this paper (see supplementary material for the sources).
@@
Authors’ psychological stress may also stem from related dynamics, as evidenced by NeurIPS submission statistics in Figure 8. Under the pressure of massive submissions, accepted papers exhibit roughly linear growth ($R^{2}=0.964,p<0.001$ under a linear regression model,), while rejections surge at a much faster, near-exponential rate ($R^{2}=0.937$, $p<0.001$ under a log-linear regression model). This inevitably drives more resubmissions, further straining reviewer capacity. Compounding the issue, state-of-the-art (SOTA) benchmarks from prior submissions may become outdated by the next cycle, fostering a distorted mindset among authors who chase overly positive reviews at the expense of genuine innovation. For instance, studies suggest that AI agent capabilities double approximately every seven months 28, 29. Since the conference cycle from submission to presentation also lasts nearly seven months, this means research can be outdated by the time it is published, rendering a significant portion of the community’s effort inefficient. Such dynamics undermine community building and knowledge sharing. Additionally, the sheer volume of acceptances burdens organizers, with exploding attendance, visa restrictions, and other logistical constraints prompting NeurIPS to adopt a hybrid online-offline format since 2022. Pushing this concept further, NeurIPS 2025 will host parallel main conferences in Mexico City 30 and Copenhagen 31. These venue-related challenges are discussed below.
@@
As academic conferences grow in scale and scope, physical venues are increasingly unable to keep pace. This pressure is particularly evident at flagship AI conferences such as NeurIPS. The Vancouver Convention Centre, host of NeurIPS 2024, has a maximum capacity of approximately 18,000 attendees 32. Recognizing that registration demand was nearing this limit (as reflected in Figure 8), NeurIPS 2024 implemented a lottery system 33 for non-author registrations. This logistical constraint has fueled community concerns that venue limits could pressure organizers to force a lower acceptance rate or even push to reject already accepted papers 34. While necessary from a logistical perspective, this policy introduced a layer of artificial scarcity that limits participation for students, early-career researchers, and unaffiliated attendees who may benefit most from in-person engagement.
@@ 5 Evaluating Incremental Fixes: Why Palliative Measures Fail
A case in point is the proposal to limit submissions per author 35 applied to manage submission volume. However, this supply-side constraint is the zero-sum game of submission caps. It does not reduce the immense institutional pressure to publish; it merely shifts it, forcing researchers to be more strategic but no less stressed. Such caps can disproportionately penalize junior researchers who need to build their publication record, or productive labs working on multiple innovative fronts. Most importantly, it does nothing to alleviate the systemic “publish-or-perish” culture that drives burnout and discourages high-risk, long-term research. It’s an administrative patch on a cultural wound.
Similarly, the adoption of multi-site or satellite conferences, such as NeurIPS 2025, hosting parallel events in Mexico City and Copenhagen 30, 31, is a direct response to venue capacity limits and aims to reduce long-haul travel for some participants. However, a multi-site conference still concentrates the entire community’s review burden and mental health anxiety into the same frantic, annual cycles. It maintains the centralized authority and the high-stakes, all-or-nothing evaluation process. It mitigates some travel but still fails to address the core issues of reviewer burnout, escalating author effort, and the exclusion inherent in any time-bound, high-cost event. It may also inadvertently create a two-tiered system, where one venue is perceived as more prestigious, thus contradicting the goal of equitable participation under the social contract.
@@ Appendix A Detailed Emission Factors
Air Travel 0.158 kg CO2e / passenger.km 36
Accommodation (Kigali, 2023) 28.0 kg CO2e / room-night 37, 19
Accommodation (Vienna, 2024) 13.9 kg CO2e / room-night 38, 39
Accommodation (Singapore, 2025) 24.5 kg CO2e / room-night 40
Catering (Meals) 1.88 kg CO2e / meal 41, 42
Venue Operations 2.896 kg CO2e / person/day 43
The data in table 2 is retrieved from the UK’s Department for Environment, Food and Rural Affairs (DEFRA) GreenHouse Gas (GHG) Conversion Factors 202544.